# Empowering Molecule Discovery for Molecule-Caption Translation with Large Language Models: A ChatGPT Perspective

### v1

@@ 3.2 Molecule-Caption Retrieval
In order to teach LLMs to handle the molecule-caption translation task (i.e., Mol2Cap and Cap2Mol) without fine-tuning LLMs, we propose to perform in-context learning with few-shot examples to prompt LLMs. In general, $n$ examples are randomly selected from human-annotated datasets (i.e., molecule-caption pair database), providing a general task instruction to LLMs. However, such a naive solution often provides insufficient knowledge regarding the associations between natural language and molecules. To mitigate this issue, we propose incorporating retrieval methods into the selection of examples to complement the lack of domain-specific knowledge of LLMs in molecule discovery, specifically through the stage of Molecule-Caption Retrieval. These retrieval strategies are motivated by the similar property principle, in which molecules similar in structures tend to exhibit similar characteristics  (Wang et al., 2016). Thus, similar captions containing the descriptions of molecule structures and properties are used to describe similar molecules. Therefore, by retrieving the most similar molecules or captions, we can utilize the corresponding molecule-caption pairs as examples to prompt LLMs.
However, the SMILES representation of molecules as a sequence structure can not reveal the actual 2-D graph topology of molecules. Hence, domain-specific methods are required for better molecular similarity calculation during the retrieval stage. Specifically, given a SMILES string representation for molecule captioning task, we introduce to use of Morgan Fingerprints (i.e., molecular structures representations) (Butina, 1999), to calculate molecular similarity using Dice similarity for molecule retrieval. In text-based molecule generation task for caption retrieval, BM25, which is widely used in information retrieval (Robertson et al., 2009), is proposed to compute similarity scores between captions of molecules, which mainly contain functional groups and properties of molecules. In both scenarios, top-n molecule-caption pair examples are retrieved to serve as examples in the system prompt. Next, we will detail Morgan Fingerprints-based molecule retrieval and BM25-based caption retrieval.
@@
3.2.1 Morgan Fingerprints-based Molecule Retrieval
Molecular fingerprints are numerical representations of the chemical structures of molecules, which can be used for various computational objectives (Butina, 1999), such as similarity searching, property prediction, virtual screening, and cluster analysis. One of the most representative molecular fingerprints is the Morgan Fingerprints (Morgan FTS), which is also known as circular fingerprints or extended-connectivity fingerprints (ECFP).
The key idea behind Morgan FTS is to capture the presence or absence of specific sub-structures or chemical fragments in a molecule. Morgan FTS follows a variant of the Morgan algorithm  (Butina, 1999), which encodes the structural information of a molecule by representing its connectivity patterns in a circular manner. Morgan FTS is then generated by iteratively expanding a set of atoms from a central atom in the molecule, capturing the neighbouring atoms and their bond types at each expansion step. The process continues until a pre-defined radius is reached. The resulting fingerprint is a binary bit vector, where each bit represents the presence or absence of a particular substructure.
What’s more, Morgan FTS has several advantages over other types of fingerprints, including their ability to handle molecules of varying sizes, resistance to small structural changes, and effectiveness in capturing structural similarities between molecules. To extract the Morgan Fingerprints, the SMILES representations of the molecules are converted into rdkit objects using the rdkit library [^1] https://www.rdkit.org/. Subsequently, we apply Dice similarity  (Dice, 1945), also known as the Dice coefficient, to measure the similarity between the input molecule and the molecules in the local database. Mathematically, it can be expressed as:
@@ 3.2.1 Morgan Fingerprints-based Molecule Retrieval
where $A$ and $B$ are the Morgan Fingerprints of the two molecules. $|A|$ and $|B|$ represent the cardinality (i.e., number of sub-structures) of $A$ and $B$. $|A\cap B|$ denotes the number of sub-structures that are common to both $A$ and $B$. Dice similarity ranges from 0 to 1, where the value of 0 indicates no overlap or similarity between the molecules, and the value of 1 represents complete overlap. As shown in Figure 3, Dice similarity can be calculated by comparing the Morgan fingerprints of the molecules. The similarity map shows the similarities and differences between the two molecules. The Dice similarity is particularly useful when dealing with imbalanced datasets or focusing on the agreement between positive instances (i.e., sub-structures present in both sets) rather than the overall agreement.
Compared to existing molecule embedding methods (Coupry & Pogány, 2022), Morgan FTS together with Dice similarity provides a distinctive advantage by explicitly indicating the similarities in detailed molecular structures, as these structures are usually directly stated in the molecule captions.
3.2.2 BM25-based Caption Retrieval
BM25 is one of the most representative ranking approaches in information retrieval for calculating the relevance of the documents to the given query  (Robertson et al., 2009). The idea is based on the term frequency-inverse document frequency (TF-IDF), which measures how often a term appears in a document (i.e., caption) and how rare it is in the corpus of documents (i.e., the local database)  (Aizawa, 2003). In addition, BM25 considers the caption’s length and the position of the query terms in the caption.
@@ 3.2.2 BM25-based Caption Retrieval
$score(Q,D)=$ $\sum_{i=1}^{N}IDF(q_{i})$
*$*\frac{f(q_{i},D)*(k_{1}+1)}{f(q_{i},D)+k_{1}*(1-b+b*\frac{|D|}{avgdl})},$ (2)
where $D$ is the caption corpus and $Q$ is the query caption. $N$ is the number of query terms in the query caption, $q_{i}$ is the $i$-th query term, $IDF(q_{i})$ is the inverse document frequency of $q_{i}$, $f(q_{i},D)$ is the term frequency of $q_{i}$ in $D$, $k_{1}$ and $b$ are tuning parameters, $|D|$ is the length of $D$, and $avgdl$ is the average caption length in the corpus.
In caption retrieval, BM25 is applied to calculate the similarity scores between captions so that the relevant molecule structures described by captions can be learnt through retrieved molecule-caption pairs.
@@ 3.3 Prompt Management
System prompts and user input prompts are two important parts to form the task context. User prompts are usually more complex and contain essential instructions for task solving and format formalization, where user prompts are defined to formalize the user inputs. To help LLMs understand the task and generate desired outputs, Prompt Management is proposed to design the system prompt templates, which are further completed with the retrieved examples. As shown in stage 2 of Figure 2, the system prompts consist of four parts: Role Identification, Task Description, Retrieved Examples, and Output Instruction.
Role Identification aims to help LLMs identify the role of experts in the chemistry and molecule discovery domain. By establishing this role, the LLMs are encouraged to generate responses that align with the expertise expected in the specific domain.
@@
The next component of the system prompt is designed to define the user input prompt and incorporate the Retrieved Examples, which serve as the evidence for the molecule-caption translation task, allowing LLMs to leverage the information contained within few-shot examples to generate better responses.
Finally, Output Instruction specifies the desired format for the response. Here, we restrict the output to a JSON format. The choice of JSON format enables a quick and efficient validation of the LLMs’ response, ensuring that it adheres to the expected structure and facilitates further processing or analysis.
@@ 3.4 In-Context Few-Shot Molecule Learning
Since ChatGPT is treated as a black-box system, it is impossible for us to fine-tune the model’s parameters on task-specific datasets for translation between molecules and natural language captions. Besides, as the weights of LLMs continue to scale, it is infeasible to train and fine-tune these foundation models with huge computational resources. To address the above limitations, recently, as an alternative to fine-tuning, in-context learning techniques provide great opportunities to teach ChatGPT to make predictions based on a few examples. In this work, we introduce in-context few-shot molecule learning to perform the Mol2Cap task and Cap2Mol task without fine-tuning ChatGPT. This stage is to utilize both the system prompt and user input prompt to query the LLMs after Molecule-Caption Retrieval and Prompt Management. In particular, the combination of the system prompt and user input prompt provides ChatGPT with a clear guideline (i.e., Mol2Cap and Cap2Mol prompts with a few examples) via in-context learning, The system prompt establishes the task framework of molecule-caption translation and molecule domain expertise, while the user prompt narrows the focus and directs the model’s attention to the specific user input. As a result, ChatGPT can learn how to perform the molecule-caption translation from the given task context, without the necessity to modify its parameters.
The formula below describes the differences between fine-tuning and in-context learning. Let $L$ be the model of ChatGPT, $m$ be the molecule, $c$ be the molecule caption, and $\theta$ be the parameters of $L$. The fine-tuning process can be formulated as:
@@
In contrast, the In-Context Few-Shot Molecule Learning process can be defined as:
@@
where $p_{m}(\cdot)$ and $p_{c}(\cdot)$ are the Prompt Management templates that transform the original user input (molecules $p_{m}$ or captions $p_{c}$) into system prompts with the user input prompts for querying ChatGPT, and $\theta$ is the original parameters without being fine-tuned.
It is apparent that the fine-tuning methods require additional model training for the sub-tasks of molecule-caption translation. In contrast, in-context few-shot molecule learning only needs to switch the prompt templates, which is much more efficient for deployment. Through the way of in-context few-shot molecule learning, valid and meaningful responses are expected, which contain the generated captions or molecules in our pre-defined JSON formats.
@@ 3.5 Generation Calibration
Despite specifying the desired output format, LLMs (e.g., ChatGPT) can occasionally produce unexpected responses, including incorrect output formats and denial of answering. To address these issues, a generation calibration mechanism is introduced to validate the response from ChatGPT.
In Generation Calibration, we first check the format of original responses by parsing them into JSON objects. If the parsing process fails, indicating a deviation from the expected format, several pre-defined format correction strategies, such as Regular Matching  (Thompson, 1968), are introduced to correct the format and extract the desired output from the response. If the original response successfully passes the format check or can be calibrated using the format correction strategies, it is considered valid and accepted as a final response. However, if the original response fails the format check and cannot be corrected within the predefined strategies, we initiate re-queries. Notably, there is a special case for re-queries. When the original response reports the ”Exceed Maximum Input Length Limitation” error, we will remove the longest example in the re-query phase until the query length meets the length limitation. The re-query process involves making additional queries to the LLMs until a valid response is obtained or until the maximum error allowance is reached. This maximum error allowance is set to ensure that the system does not get stuck in an endless loop and instead provides a suitable response to the user within acceptable bounds.
By employing the generation calibration stage, we can mitigate unexpected deviations from the desired output format and ensure that the final responses align with the expected format and requirements.
@@ 4 Experiment
In this section, we aim to evaluate the feasibility and effectiveness of the proposed method MolReGPT by conducting comprehensive experiments on the molecule-caption translation task. Additionally, ablation studies are conducted to investigate the impact of different retrieval methods and the number of selected examples. These investigations aim to provide deeper insights into the performance and capabilities of MolReGPT.
@@ 4.1 Experimental Settings
We first introduce the basic experimental settings. In this work, we use ChatGPT through the OpenAI API  [^2] https://openai.com/blog/openai-api with backend model GPT-3.5-turbo, which can not be fine-tuned in our tasks. Besides, we will provide an overview of the data and metrics employed in this section.
4.1.1 Dataset
@@
The research on molecule-caption translation is still in the early stage, and there is only one public dataset ChEBI-20 (Edwards et al., 2021), which contains 33,010 molecule-caption pairs. To ensure consistency, we adhere to the data split process as used in MolT5  (Edwards et al., 2022), dividing the dataset into 80/10/10% train/validation/test splits. For our method evaluation, we focus on the test split while utilizing the training set as the local database to retrieve n-shot examples through in-context learning.
4.1.2 Evaluation Metrics
@@
In terms of evaluation metrics, we align with the metrics adopted in MolT5 for comparison  (Edwards et al., 2022). By adopting these metrics, we ensure consistency and enable a meaningful and fair assessment of the performance of our method.
  • Mol2Cap Metrics. In the Mol2Cap task, natural language generation metrics like BLEU, ROUGE, and METEOR scores are applied to assess the proximity of the generated output to the ground truth. Here, BLEU and ROUGE scores evaluate the n-gram precision, measuring the alignment between the generated structures and the reference structures, while METEOR is a recall-oriented metric that accounts for both exact matches and paraphrases between the generated and reference structures. Additionally, we incorporate Text2Mol, a task-specific metric that employs pre-trained models to quantify the structural similarity between the generated and reference molecules using their SMILES representations (Edwards et al., 2021). This metric provides further insights into the quality and relevance of the generated output in terms of the underlying molecular structures.
  • Cap2Mol Metrics. Since the SMILES representation of molecules exhibits a sequence structure, natural language metrics can be directly applied for evaluation. Thus, BLEU and the Exact Match scores are calculated as initial assessments. Furthermore, molecule-specified metrics are also reported, including Levenshtein distance, validity, and three molecule fingerprints scores - MACCS FTS, RDK FTS, and Morgan FTS. These metrics provide valuable insights into the quality, validity, and structural characteristics of the generated molecules. Finally, the Text2Mol metric is also discussed here to highlight the relevance between the generated molecules and the input molecule captions.
Note that smaller values of Levenshtein score and FCD indicate better generation performance in the molecule generation task, while other evaluation metrics positively correlate to the performance.
4.1.3 Baselines
It is worth mentioning that there are limited baselines for translating between molecule captioning and text-based molecule generation. Specifically, the following baselines are selected for performance evaluation:
@@ 4.1.3 Baselines
  • Transformer (Vaswani et al., 2017). This method is the most representative language architecture to process natural language. A vanilla Transformer model with six encoder and decoder layers, directly trained on ChEBI-20. Note that this model is not pre-trained, making it simple and easy to implement.
@@
  • T5-base (Raffel et al., 2020). T5 is pre-trained on the Colossal Clean Crawled Corpus (C4), but no domain knowledge is specifically fed for pre-training. In this work, the base version of T5 is directly fine-tuned on ChEBI-20 for molecule discovery.
@@
  • MolT5-base (Edwards et al., 2022). This model is pre-trained on a large corpus with both language texts and SMILES strings so that it can have a prior understanding of the two domains. More specifically, the base version of MolT5 was pre-trained on the Colossal Clean Crawled Corpus (C4) and ZINC-15 datasets and further fine-tuned on task-specific dataset ChEBI-20.
Note that these baselines are required to fine-tune the model on the public dataset ChEBI-20, specifically tailored to the molecule-caption translation task.
4.2 Performance Comparison of Molecule-Caption Translation

### v2

@@ 3.1 Molecule-Caption Retrieval
In order to teach LLMs to handle the molecule-caption translation task without fine-tuning LLMs, we propose in-context few-shot molecule learning to guide LLMs to learn how to translate between molecules and text captions. Normally, in-context learning requires $n$ random examples selected from human-annotated datasets (i.e., molecule-caption pair database), providing a general task instruction to LLMs. However, random examples often provide insufficient knowledge regarding the associations between natural language and molecules, as they fail to provide useful information for the detailed descriptions of functional groups and molecule features. To mitigate this issue, we propose incorporating retrieval methods into the selection of context examples to complement the lack of task-specific knowledge in LLMs, specifically through the stage of Molecule-Caption Retrieval. These retrieval strategies are motivated by the similar property principle, in which molecules similar in structures tend to exhibit similar characteristics  [37]. Namely, similar captions containing the descriptions of molecule structures and properties are used to describe similar molecules. Therefore, via these most similar molecules or captions, we could utilize the corresponding molecule-caption pairs as context examples to guide LLMs.
Notably, the SMILES representation of molecules, as a sequence structure, can hardly reveal the actual 2-D graph topology of molecules. Hence, domain-specific methods are required for better molecular similarity calculation during the retrieval stage. Specifically, given a SMILES string representation for Mol2Cap task, we introduce using Morgan Fingerprints (i.e., a molecular structures representation) [38] to calculate molecular similarity using Dice similarity for molecule retrieval. Meanwhile, in Cap2Mol task, we are more focused on the details in the text captions (e.g., IUPAC names and functional group positions). Thus, BM25 caption retrieval, which is widely used in information retrieval [39], is proposed to compute similarity scores between captions of molecules.In both sub-tasks, top-n molecule-caption pair examples are retrieved to serve as context examples in the system prompt. Next, we will detail Morgan Fingerprints-based molecule retrieval and BM25-based caption retrieval.
3.1.1 Morgan Fingerprints-based Molecule Retrieval.
Molecular fingerprints are numerical representations of the chemical structures of molecules, which can be used for various computational objectives [38], such as similarity searching, property prediction, and cluster analysis. One of the most representative molecular fingerprints is the Morgan Fingerprints (Morgan FTS).
@@
The key idea behind Morgan FTS is to capture the presence or absence of specific sub-structures or chemical fragments in a molecule. Morgan FTS follows a variant of the Morgan algorithm [38], which encodes the structural information of a molecule by representing its connectivity patterns in a circular manner. Morgan FTS is then generated by iteratively expanding a set of atoms from a central atom in the molecule, capturing the neighbouring atoms and their bond types at each expansion step. The process continues until a pre-defined radius is reached. The result is a binary bit vector, where each bit represents the presence or absence of a particular substructure.
What’s more, Morgan FTS has several advantages over other types of fingerprints, including their ability to handle molecules of varying sizes, resistance to small structural changes, and effectiveness in capturing structural similarities between molecules. Subsequently, we apply Dice similarity [40] to measure the similarity between the input molecule and the molecules in the local database. Mathematically, it can be expressed as:
@@
where $A$ and $B$ are the Morgan Fingerprints of two molecules. $|A|$ and $|B|$ represent the cardinality (i.e., number of sub-structures) of $A$ and $B$. $|A\cap B|$ denotes the number of sub-structures that are common to both $A$ and $B$. Dice similarity ranges from 0 to 1. The Dice similarity is particularly useful when dealing with imbalanced datasets or focusing on the agreement between positive instances (i.e., sub-structures present in both sets) rather than the overall agreement. As shown in Figure 3, the similarity map explains how Morgan FTS measures the similarities and differences between the two molecules.
Compared to existing molecule embedding methods, Morgan FTS together with Dice similarity provides a distinctive advantage by explicitly indicating the similarities in detailed molecular structures, as these structures are usually directly stated in the molecule captions [41].
3.1.2 BM25-based Caption Retrieval.
BM25 is one of the most representative ranking approaches in information retrieval for calculating the relevance of the documents to the given query. The idea is based on the term frequency-inverse document frequency (TF-IDF), which measures how often a term appears in a document (i.e., caption) and how rare it is in the corpus of documents (i.e., the local database)  [42]. In addition, BM25 considers the caption’s length and the position of the query terms in the caption.
@@
$score(Q,\!D)\!=\!\sum_{i=1}^{N}\!I\!D\!F(\!q_{i}\!)\!*\!\frac{f(q_{i},D)*(k_{1}+1)}{f(\!q_{i},\!D\!)\!+\!k_{1}\!*\!(\!1\!-\!b\!+\!b\!*\!\frac{|D|}{avgdl}\!)},$ (2)
where $D$ is the caption corpus and $Q$ is the query caption. $N$ is the number of query terms in the query caption, $q_{i}$ is the $i$-th query term, $IDF(q_{i})$ is the inverse document frequency of $q_{i}$, $f(q_{i},D)$ is the term frequency of $q_{i}$ in $D$, $k_{1}$ and $b$ are tuning parameters, $|D|$ is the length of $D$, and $avgdl$ is the average caption length in the corpus. In Cap2Mol task, as we discussed, structure details usually bring more information gain to the prediction of the molecule SMILES representations, while such details are difficult to capture at the semantic level. In this case, BM25 caption retrieval is applied to calculate the similarity scores between captions at the token level so that the relevant molecule structures described by captions can be learnt via molecule-caption pairs.
@@ 3.2 Prompt Management
System prompts and user input prompts are two important parts to enable the in-context learning ability of LLMs. User prompts are usually more complex and contain essential instructions for task solving and format formalization, where user prompts are defined to formalize the user inputs. To help LLMs understand the task and generate desired outputs, Prompt Management is proposed to design the system prompt templates, which are further completed with the context examples. As shown in stage 2 of Figure 2, the system prompts consist of the following four parts:
  • Role Identification aims to help LLMs identify the role of experts in the chemistry and molecule discovery domain. By establishing this role, LLMs are encouraged to generate responses aligning with the expected expertise.
@@
  • Context Examples serves as the evidence for the molecule-caption translation task, allowing LLMs to leverage the information contained within the molecule-caption pairs via in-context learning to generate better responses.
  • Output Instruction specifies the desired format for the response. Here, we restrict the output to a JSON format for further processing or analysis.
@@ 3.3 In-Context Few-Shot Molecule Learning
Recently, as an alternative to fine-tuning, in-context learning provides great opportunities to teach LLMs to make predictions based on a few context examples. In this work, we introduce In-Context Few-Shot Molecule Learning to perform the Mol2Cap and Cap2Mol tasks without fine-tuning LLMs and analyse the underlying mechanism. This stage is to utilize both the system prompt and user input prompt to query the LLMs. In particular, the combination of the system prompt and user input prompt provides LLMs with a clear guideline via in-context learning; the system prompt establishes domain expertise for molecule-caption translation, while the user prompt narrows the focus and directs the model’s attention to the specific user input. As a result, LLMs can learn how to perform the molecule-caption translation from the given task context, without the necessity to modify their parameters.
The formulas below describe the differences between fine-tuning, prompting, and In-context Few-Shot Molecule Learning. Let $M$ be the large language model, $m$ be the molecule, $c$ be the corresponding molecule caption, and $\theta$ be the parameters of the LLM.
Fine-tuning process for the Mol2Cap and Cap2Mol task aims to help LLMs learn the translation probability between the molecule and its text caption:
@@
Prompting uses prompts to wrap the inputs and guide the generation of LLMs without changing the parameters. It can be defined as:
@@
where $p_{m}(\cdot)$ and $p_{c}(\cdot)$ are the Prompt Management templates that transform the original user input (molecules $p_{m}$ or captions $p_{c}$) into system prompts with the user input prompts for querying LLMs, and $\theta$ is the original parameters without being fine-tuned.
In-Context Few-Shot Molecule Learning targets the alignment between text and molecule structures. Our method can be formulated as:
$c=L(p_{m}(m)||T_{m\rightarrow c};\theta),$ (7)
$m=L(p_{c}(c)||T_{c\rightarrow m};\theta),$ (8)
where $T_{m\rightarrow c}=(m_{1}\rightarrow c_{1})||(m_{2}\rightarrow c_{2})||...||(m_{n}\rightarrow c_{n})$ and $T_{c\rightarrow m}=(c_{1}\rightarrow m_{1})||(c_{2}\rightarrow m_{2})||...||(c_{n}\rightarrow m_{n})$ are the retrieved context examples for the Mol2Cap and Cap2Mol task. Here, $(c_{i}\rightarrow m_{i})$ denotes the $i$-th retrieved example that illustrates how to transfer from caption $c_{i}$ to the molecule $m_{i}$, while $(m_{i}\rightarrow c_{i})$ represents the alignment from the molecule $m_{i}$ to its caption $c_{i}$, where $i\in[1,n]$. Notably, $\theta$ is still the original parameter of LLMs without any modification.
Through the way of learning the molecule-caption translation from context examples, LLMs can grasp the alignment between text description and the molecule structures it implies. Compared to fine-tuning, our method does not require additional model training. Compared to prompting, our method is more explainable and robust due to the in-context learning process.
@@ 3.4 Generation Calibration
Despite specifying the desired output format, LLMs (e.g., ChatGPT) can occasionally produce unexpected responses, including incorrect output formats and denial of answering. To address these issues, a generation calibration mechanism is introduced to validate the response from LLMs.
In Generation Calibration, we first check the format of original responses. If the format is not correct, several pre-defined format correction strategies, such as Regular Matching, are introduced to correct the format and extract the desired output from the response. If the original response passes the format check or can be calibrated, it is considered valid and accepted as a final response. However, if the original response fails the format check and cannot be corrected within the predefined strategies, we initiate re-queries. Notably, there is a special case for re-queries. When the original response reports the “Exceed Maximum Input Length Limitation” error, we will remove the longest example in the re-query phase until the query length meets the length limitation. The re-query process involves making additional queries to the LLMs until a valid response is obtained or until the maximum error allowance is reached to ensure that the system does not get stuck in an endless loop.
@@ 4 Experiment
[…]
Transformer [10] 0.061 0.027 0.204 0.087 0.186 0.114 0.057
GPT-3.5-turbo (zero-shot) 0.103 0.050 0.261 0.088 0.204 0.161 0.352
LLama-2-7B (zero-shot) 0.094 0.039 0.169 0.054 0.142 0.175 0.153
LLama-2-7B (2-shot MolReGPT) 0.489 0.409 0.535 0.374 0.472 0.495 0.466
T5-base [10] 0.511 0.423 0.607 0.451 0.550 0.539 0.523
MolT5-base [10] 0.540 0.457 0.634 0.485 0.578 0.569 0.547
T5-large [10] 0.558 0.467 0.630 0.478 0.569 0.586 0.563
MolT5-large [10] 0.594 0.508 0.654 0.510 0.594 0.614 0.582
GPT-4-0314 (10-shot MolReGPT) 0.607 0.525 0.634 0.476 0.562 0.610 0.585
TABLE I: The performance of Mol2Cap on ChEBI-20. Experimental results for Transformer, T5-base, MolT5-base, T5-large, and MolT5-large are retrieved from  [10]. Due to the input length limitation, we apply 2-shot MolReGPT to llama-2-7B and 10-shot MolReGPT to GPT models. The best scores are in bold, and the second-best scores are underlined.
[…]
Transformer [10] 0.499 0.000 57.66 0.480 0.320 0.217 11.32 0.277 0.906
LLama-2-7B (zero-shot) 0.104 0.000 84.18 0.243 0.119 0.089 42.01 0.148 0.631
LLama-2-7B (2-shot MolReGPT) 0.693 0.022 36.77 0.808 0.717 0.609 4.90 0.149 0.761
T5-base [10] 0.762 0.069 24.950 0.731 0.605 0.545 2.48 0.499 0.660
MolT5-base [10] 0.769 0.081 24.458 0.721 0.588 0.529 2.18 0.496 0.772
T5-large [10] 0.854 0.279 16.721 0.823 0.731 0.670 1.22 0.552 0.902
MolT5-large [10] 0.854 0.311 16.071 0.834 0.746 0.684 1.20 0.554 0.905
GPT-4-0314 (10-shot MolReGPT) 0.857 0.280 17.14 0.903 0.805 0.739 0.41 0.593 0.899
TABLE II: Cap2Mol results on ChEBI-20. Experimental results for Transformer, T5-base, MolT5-base, T5-large, and MolT5-large are retrieved from [10]. Due to the input length limitation, we apply 2-shot MolReGPT to llama-2-7B and 10-shot MolReGPT to GPT models. The best scores are in bold, and the second-best scores are underlined.
In this section, we aim to evaluate the effectiveness of our proposed MolReGPT by conducting comprehensive experiments on molecule-caption translation task.
@@ 4.1 Experimental Settings
We first introduce the basic experimental settings. In this work, we use ChatGPT through the OpenAI API [^1] https://openai.com/blog/openai-api with backend model GPT-3.5-turbo and GPT-4-0314, which can not be fine-tuned in our tasks. To assure the model agnosticism of MolReGPT, we also apply an extra open-source LLM, Llama-2-7b-chat-hf (i.e., Llama-2-7B) [43], for comparison. Notably, due to the input length limitation and hardware limitation, we could only apply 2-shot MolReGPT to this model. Besides, we will provide an overview of the data and metrics employed in this section.
@@
Dataset. The research on molecule-caption translation is still in the early stage, and there is only one public dataset ChEBI-20 [9], which contains 33,010 molecule-caption pairs. To ensure consistency, we adhere to the data split process as used in MolT5  [10], dividing the dataset into 80/10/10% train/validation/test splits. For our method evaluation, we focus on the test split while utilizing the training set as the local database to retrieve n-shot examples for in-context learning.
@@
Evaluation Metrics. In terms of evaluation metrics, we align with the metrics adopted in MolT5  [10]. By adopting these metrics, we ensure consistency and enable a fair assessment of the performance of our method.
Baselines. Specifically, the following baselines are selected for performance evaluation. Most of these baselines are only fine-tuned on ChEBI-20 because of their limited maximum input length and poor reasoning capabilities.
@@
  • Transformer [44]. This method is the most representative language architecture to process natural language. A vanilla Transformer model with six encoder and decoder layers, directly trained on ChEBI-20. Notably, this model is not pre-trained.
@@
  • T5 [45]. T5 is pre-trained on the Colossal Clean Crawled Corpus (C4), but no domain knowledge is specifically fed for pre-training. In this work, T5-base and T5-large are directly fine-tuned on ChEBI-20.
@@
  • MolT5 [10]. MolT5 models are pre-trained on both language texts and SMILES strings. More specifically, MolT5-base and MolT5-large were pre-trained on the Colossal Clean Crawled Corpus (C4) and ZINC-15 datasets and further fine-tuned on ChEBI-20.
4.2 Performance Comparison